solve the system of equations by equations by elimanation
9x+5y=11
4x+y=11

Answer by checkley79(3341)

9x+5y=11
4x+y=11 MULTIPLY BY -5 & ADD.
-20X-5Y=-55
9X+5Y=11
-------------------
-11X=-44
X=-44/-11
X=4 ANS.
9*4+5Y=11
36+5Y=11
5Y=11-36
5Y=-25
Y=-25/5
Y=-5 ANS.
PROOF:
4*4+(-5)=11
16-5=11
11=11